OpenStudy (anonymous):

so if it was t^8 (t^10)^0 thats t^8and t^10 I would add 8 and 10 so its t^18? am i right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its t^80

OpenStudy (anonymous):

any number to the power of zero is 1

OpenStudy (anonymous):

\((t^8)^0=1\)

OpenStudy (anonymous):

1x10=10 and 10x8= 80?

OpenStudy (anonymous):

you are raising \(t^8\) to the power of zero, not to the power of one

OpenStudy (anonymous):

\((t^8)^0=t^{8\times 0}=t^0=1\)

OpenStudy (anonymous):

thats for the 10 i think

OpenStudy (anonymous):

in general \(b^0=1\) for any number \(b\) (unless \(b=0\) but that is a different story)

OpenStudy (anonymous):

my fault, but same idea

OpenStudy (anonymous):

\((t^{10})^0=t^0=1\)

OpenStudy (anonymous):

so its just t^8?

OpenStudy (anonymous):

to you are left with \[t^8\times t^0=t^8\times 1=t^8\]

OpenStudy (anonymous):

yes, it is \(t^8\) you are right

OpenStudy (whpalmer4):

Think of it like t^2 = t*t, so t^1 = t*t/t = t and t^0 = t*t/(t*t) = 1

OpenStudy (anonymous):

thxs you all